# Euler characteristic of crepant resolutions of specific modular quotient singularities

### v1

@@ Abstract
In this paper, we consider a generalization of McKay correspondence in positive characteristic regarding Euler characteristic of crepant resolutions given by finite subgroups of the special linear group. As the main result, we prove that the generalization holds for groups with a specific semidirect product structure. Apart from such groups, two certain examples for groups with more difficult structure are also given.
@@ 1 Introduction
Euler characteristic of crepant resolutions is one of important geometric invariants in study of McKay correspondence. As a series of results connecting algebra with geometry, McKay correspondence studies relations between algebraic properties of groups and geometric properties of the associated quotient singularities. Over complex numbers, it was conjectured by Reid and proved by Batyrev via motivic integration ([batyrev1999non]), that for a finite group $G\subseteq\mathrm{SL}(n,\mathbb{C})$ and the associated quotient singularity $X:=\mathbb{C}^{n}/G$, assuming existence of a crepant resolution $Y\to X$, the Euler characteristic $e(Y)$ equals the number of conjugacy classes #$\#\mathrm{Conj}(G)$.
In this paper, the case when the base field is of positive characteristic is focused. Let $k$ be an algebraically closed field of prime characteristic $p>0$, and $G\subseteq\mathrm{SL}(n,k)$ be a finite group giving the quotient singularity $X:=\mathbb{A}^{n}_{k}/G$. For a smooth $k$-variety, its (topological) Euler characteristic is defined as the alternating sum of Betti numbers given by $l$-adic cohomology with compact support, where $l$ is a prime different from $p$.
It is natural to ask if the equality between Euler characteristic of crepant resolutions and number of conjugacy classes still holds. If the order of $G$ is not divided by $p$, then there is an affirmative answer known as tame McKay correspondence (the reader can refer to [wood2015mass], Section 5.2). Unfortunately, in modular case, where #$\#G$ is divided by $p$, the analog of Batyrev’s theorem has been disproved by construction of counterexamples, such as Chen, Du and Gao’s counterexample in non-small case ([chen2020modular]), Yamamoto’s construction for the quotient variety given by the symmetric group $S_{3}\subseteq\mathrm{SL}(3)$ in characteristic $3$ ([yamamoto2021crepant]) and the author’s construction for the quotient variety given by permutation action of the alternating group $A_{4}$ in characteristic $2$ ([fan2023crepant]).
On the other hand, there are also several known results implying that the analog of Batyrev’s theorem in positive characteristic holds for specific modular quotient singularities. We list them as below.
@@
    ([yasuda2014cyclic], Corollary 6.21) Let $k$ be an algebraically closed field of characteristic $p>0$. Suppose $G\subseteq\mathrm{SL}(n,k)$ is a finite group with no reflections, such that $G\cong C_{p}$, where $C_{p}$ is the $p$-cyclic group. If the associated quotient singularity has a crepant resolution $Y\to X:=\mathbb{A}^{n}_{k}/G$, then $e(Y)=p$.
@@
    ([yamamoto2021crepant], Theorem 1.2) Let $k$ be an algebraically closed field of characteristic $3$. Suppose $G\subseteq\mathrm{SL}(3,k)$ is a finite group with no reflections, such that $G\cong H\rtimes C_{3}$, where $H$ is a non-modular abelian group and $C_{3}$ is the cyclic group of order $3$. Then the associated quotient singularity has a crepant resolution $Y\to X:=\mathbb{A}^{3}_{k}/G$, and $e(Y)=\#\mathrm{Conj}(G)$.
@@
    ([fanmaster], Corollary 3.1) Let $k$ be an algebraically closed field of characteristic $p=2$. For any positive integer $l$ with $(p,l)=1$, let $\zeta_{l}$ be a primitive $l$-th root of unity in $k$. For integers $a_{1},a_{2},a_{3},a_{4}$, denote the diagonal matrix $\mathrm{diag}(\zeta_{l}^{a_{1}},\zeta_{l}^{a_{2}},\zeta_{l}^{a_{3}},\zeta_{l}^{a_{4}})$ by $\frac{1}{l}(a_{1},a_{2},a_{3},a_{4})$. Suppose $G\subseteq\mathrm{SL}(4,k)$ has a semidirect product structure $G\cong H\rtimes C_{2}$, where $C_{2}$ is generated by the permutation $(12)(34)$ as an element in $\mathrm{SL}(4)$, and $H$ is one of the following:
@@
    Then $G$ has no reflections, and the associated quotient singularity has a crepant resolution $Y\to X:=\mathbb{A}^{4}_{k}/G$, satisfying $e(Y)=\#\mathrm{Conj}(G)$.
In the results above, $G$ is always small (that is, $G$ has no reflections) and has a semidirect product structure given by a non-modular abelian normal subgroup and a $p$-cyclic group. If $G\cong H\rtimes K$, it is a classical idea to consider the decomposition $\mathbb{A}^{n}\xrightarrow{/H}\mathbb{A}^{n}/H\xrightarrow{/K}\mathbb{A}^{n}/G$ and the singularities given by $H$ and $K$ respectively. Under the assumption, $H$ is non-modular, and the singularity given by $K\cong C_{p}$ is studied by Theorem 1.1.1, with the same statement as Batyrev’s theorem. In fact, Theorem 1.1.2 and Theorem 1.1.3 are parallel to Ito’s construction of crepant resolutions of trihedral singularities ([ito1994crepant]) and specific $4$-dimensional quotient singularities ([hayashi2017existence]) over $\mathbb{C}$, with the same approach. In this paper, we show that for such small groups with semidirect product structures, the analogous statement of Batyrev’s theorem always holds.
@@
Let $k=\overline{\mathbb{F}_{p}}$ be the algebraic closure of the finite fields of characteristic $p>0$, and $G\subseteq\mathrm{SL}(n,k)$ be a finite group with no reflections, such that $G$ has a structure of semidirect product as $H\rtimes C_{p}$, where $H$ is a non-modular abelian group and $C_{p}$ denotes the $p$-cyclic group. If $X:=\mathbb{A}^{n}_{k}/G$ has a crepant resolution $f:Y\to X$, then
[…]
Hereby, Theorem 1.1.2 and Theorem 1.1.3 become examples of the main theorem. By this result, we also point out that the number of isomorphic classes of indecomposable $kG$-modules, which is not necessarily equal to #$\#\mathrm{Conj}(G)$ in positive characteristic, is a candidate for the algebraic invariant of $G$ corresponding to Euler characteristic of the crepant resolution. Indeed, we have the following conjecture.
@@
Let $k$ be an algebraically closed field of characteristic $p>0$, and $G\subseteq\mathrm{SL}(n,k)$ be a finite small group of finite representation type (that is, there are finitely many isomorphic classes of indecomposable $kG$-modules). If $X:=\mathbb{A}^{n}_{k}/G$ has a crepant resolution $f:Y\to X$, then
[…]
This conjecture is true for all known cases, including non-modular case, the main theorem and Yamamoto’s counterexample in [yamamoto2021crepant]. On the other hand, there are also examples where the group is of infinite representation type, such as the author’s construction in [fan2023crepant]. In the last section, two quotient singularities given by certain groups of infinite representation type are considered.
@@
Let $k$ be an algebraically closed field of characteristic $2$, $A_{4}$ and $C_{2}^{2}$ be subgroups of $\mathrm{SL}(4,k)$ given by permutation actions of the alternating group and its normal subgroup containing elements of order $2$ respectively. Denote the associated quotient singularities by $X_{1}:=\mathbb{A}^{4}/{C_{2}^{2}}$ and $X_{2}:=\mathbb{A}^{4}/A_{4}$. If there exists a crepant resolution $Y_{1}\to X_{1}$ ($resp.$ $Y_{2}\to X_{2}$), then $e(Y_{1})=6$ ($resp.$ $e(Y_{2})=10$).
This paper is organized as follows. In Section 2, we introduce results and notations used for the proof of the main theorem, including the main tool - mass formula version of wild McKay correspondence. In Section 3, we prove the main theorem when $G$ is abelian. In Section 4, we prove the main theorem when $G$ is non-abelian. In Section 5, we provide two examples of computation of Euler characteristic of crepant resolutions for groups of infinite representation type, using key idea of proof of the main theorem.
Acknowledgements.
The author expresses his gratitude to Prof. Takehiko Yasuda and Dr. Rin Gotou for valuable discussions. The author also thanks his supervisor, Prof. Yukari Ito, for her helpful advice and consideration.
@@ 2 Preliminaries
For the proof of our main theorem, wild McKay correspondence as mass formulas is an important tool. In this version of motivic wild McKay correspondence over finite fields, stringy motives are studied by its realization as stringy point-count #$\#_{\mathrm{st}}$. Under specific conditions, if crepant resolutions exist, then some of their geometric properties can be studied via stringy point-count of the quotient varieties. For more details, the reader can refer to [yasuda2017wild]. Here below we only list results that are necessary for our proof.
@@
Let $K$ be a field and $G$ be a finite group. $M$ is called a $G$-étale $K$-algebra, if $M$ is a finite étale $K$-algebra of degree #$\#G$ equipped with a $G$-action, such that $M^{G}=K$. Isomorphisms of $G$-étale $K$-algebras are isomorphisms of $K$-algebras that are $G$-equivariant. $G\text{-}\mathrm{\acute{E}t}(K)$ is the set consisting of isomorphic classes of $G$-étale $K$-algebras.
@@
Different actions of $G$ on the same $K$-algebra may result in different $G$-étale $K$-algebras. However, conjugate endomorphisms of $G$ do not change the class in $G\text{-}\mathrm{\acute{E}t}(K)$. For each $M\in G\text{-}\mathrm{\acute{E}t}(K)$, there exists a subgroup $G_{M}\subseteq G$ (unique up to conjugation) such that $M\cong L_{M}^{\oplus(G:G_{M})}$, where $L_{M}$ is a Galois extension of $K$ with Galois group isomorphic to $G_{M}$.
Theorem 2.3 ([yasuda2017wild], Proposition 8.5).
Let $k$ be a finite field, $\mathbb{F}_{q}/k$ be a field extension with order $q=p^{e}$, and $K=\mathbb{F}_{q}((t))$ be the local field of formal Laurent series. Let $G$ be a finite subgroup of $\mathrm{SL}(n,k)$ with no reflections. Let $X:=\mathbb{A}^{n}/G$ be the associated quotient variety. Then
#$\#_{\mathrm{st},\mathbb{F}_{q}}X=\sum_{M\in G\text{-}\mathrm{\acute{E}t}(K)}\frac{q^{n-\mathbf{v}_{V}(M)}}{\#C_{G}(G_{M})};$
@@
#$\#Y(\mathbb{F}_{q})=\#_{\mathrm{st},\mathbb{F}_{q}}X=\sum_{M\in G\text{-}\mathrm{\acute{E}t}(K)}\frac{q^{n-\mathbf{v}_{V}(M)}}{\#C_{G}(G_{M})}.$
Here #$\#Y(\mathbb{F}_{q})$ counts number of $\mathbb{F}_{q}$-rational points of $Y$, $C_{G}(G_{M})$ is the centralizer of $G_{M}$ in $G$, and $\mathbf{v}_{V}(M)$ is the $\mathbf{v}$ function defined as follows.
Definition 2.4 ([yasuda2017wild]; $\mathbf{v}$ function).
$\mathbf{v}$ function is defined for an arbitrary finite group $G$ with a representation $\rho:G\to\mathrm{GL}(n,k)\hookrightarrow\mathrm{GL}(n,\mathcal{O}_{K})$. Let $V=\mathbb{A}^{n}_{\mathcal{O}_{K}}$ be the representation space and $\mathcal{O}_{K}[x_{1},\dots,x_{n}]$ be its coordinate ring with linear part $T=\sum\mathcal{O}_{K}x_{i}$. Then for any $M\in G\text{-}\mathrm{\acute{E}t}(K)$, $G$ has both an action on $\mathcal{O}_{M}$ via definition of $G$-étale $K$-algebras and an action on $T$ via the representation $\rho$. Denote by $\mathrm{Hom}_{\mathcal{O}_{K}}^{G}(T,\mathcal{O}_{M})$ the homomorphisms that commute with the actions of $G$. Then $\mathrm{Hom}_{\mathcal{O}_{K}}^{G}(T,\mathcal{O}_{M})$ is a free $\mathcal{O}_{K}$-module of rank $n$ ([yasuda2017toward], Proposition 6.3), and
@@
$\mathbf{v}$ function is convertible and additive ([wood2015mass], Lemma 3.4). We would use following facts: Even if $\rho$ is not faithful, it can be decomposed as $G\rightarrow G/H\rightarrow\mathrm{GL}(n,k)$ for some $H\subseteq G$, and $\mathbf{v}_{V}(M)=\mathbf{v}_{V_{G/H}}(M^{H})$; if $V$ can be decomposed as $V=\bigoplus_{i=1}^{s}V_{i}$, then $\mathbf{v}_{V}=\sum_{i=1}^{s}\mathbf{v}_{V_{i}}$.
@@
The value of $\mathbf{v}$ function for $M\in G\text{-}\mathrm{\acute{E}t}(K)$ is determined by $L_{M}/K$ as a $G_{M}$-étale $K$-algebra. We simply write $\mathbf{v}(L)$ if $G_{M}$ is fixed.
Next, we give several notations and use them to rewrite the wild McKay correspondence such that Euler characteristic of crepant resolutions can be studied.
@@
$F_{G}(q):=\sum_{[G^{\prime}]\subseteq G}\frac{1}{\#N_{G}(G^{\prime})}f_{G^{\prime}}.$
Here $N_{G}(G^{\prime})$ is the normalizer of $G^{\prime}$ in $G$, and $[G^{\prime}]\subseteq G$ is taken up to conjugation.
@@
Every element in $A(G)$ contains information of a Galois extension $L/K$ with Galois group $\mathrm{Gal}(L/K)\cong G$ and the action of $G$ on it. For example, if $G\cong C_{l}$ for some prime $l$ ($l\neq p$), then by Kummer theory, there are $l+1$ Galois extensions of $K=k((t))$ with $l$-cyclic Galois group up to isomorphism: one is unramified and the other $l$ extensions are totally ramified. On the other hand, for each extension, there are $l-1$ ways to map the given generator of $G$ to a nontrivial element in the Galois group, giving different actions up to isomorphism of $G$-étale algebras. Hence $A(C_{l})$ has $l^{2}-1$ elements.
@@
Let $f$ be a Laurent polynomial with coefficients in $\mathbb{Q}$. We define $S(f):=f(1)$ as the sum of coefficients of $f$. $S$ is therefore a homomorphism between $\mathbb{Q}$-algebras.
@@
Under the assumption of Theorem 2.3, if crepant resolution $Y$ exists, then $F_{G}(T^{2})$ converges as a $\mathbb{Z}$-coefficient polynomial in $T$, and the (topological) Euler characteristic $e(Y)=S(F_{G})$.
@@
The first assertion is from motivic wild McKay correspondence (construction of stringy motives and their realization as Poincaré polynomials of crepant resolutions via $\mathbb{L}\mapsto T^{2}$), and the second one follows; one can also check it by the functional equation in Weil conjectures. The reader can refer to [yasuda2014cyclic] for details. ∎
@@
If $G$ is non-modular, then $F_{G}$ is a polynomial in $q$ and $S(F_{G})=\#\mathrm{Conj}(G)$. If $G$ is also abelian, then $\sum_{G^{\prime}\subseteq G}f_{G^{\prime}}=(\#G)^{2}$.
@@
This is a corollary of tame McKay correspondence ([wood2015mass], Section 5.2). We also point out that $S(f_{G})=\#A(G)$ here. ∎
@@
If $G$ is isomorphic to a cyclic group of order $p$ and there exists a crepant resolution $Y\to X$ , then $F_{G}$ is a polynomial in $q$, $S(F_{G})=p$ and $S(f_{G})=p^{2}-1$.
@@
This is a corollary of $p$-cyclic McKay correspondence ([yasuda2014cyclic], Corollary 6.21). ∎
In the last part of this section, we give the proof for the equality between number of conjugacy classes and number of isomorphic classes of indecomposable representations in the statement of main theorem.
@@
Let $G$ be a finite group equipped with a semidirect product structure $H\rtimes C_{p}$, where $H$ is abelian and $C_{p}$ is a cyclic group of prime order $p$. Let $C(G)$ be the center of $G$ and $k$ be a splitting field of $G$ of characteristic $p$ (equivalently, for any element $h\in H$, $k$ contains all of the $\mathrm{ord}(h)$-th primitive roots of unity in $\overline{k}$). Then
#$\#\mathrm{Conj}(G)=\left\{\begin{array}[]{ll}\displaystyle p\#C(G)+\frac{\#H-\#C(G)}{p}&G\text{ is not abelian,}\\ \#G&G\text{ is abelian.}\end{array}\right.$
which coincide the number of isomorphic classes of indecomposable $kG$-modules.

### v2

@@ Abstract
In this paper, we consider a generalization of the McKay correspondence in positive characteristic regarding the Euler characteristic of crepant resolutions of quotient singularities given by finite subgroups of the special linear group. As the main result, we prove that this generalization holds for groups with a specific semidirect product structure, using the wild McKay correspondence over finite fields as mass formulas. Furthermore, two additional examples with more complicated structures are also given. Based on our main result, we propose a conjectural form of the generalized McKay correspondence in the modular case.
keywords:
McKay correspondence , crepant resolution , Euler characteristic , positive characteristic , modular representation theory , local Galois representations
2020 MSC:
14E16 , 20C20 , 11S15
journal: Journal of Algebra
[…]
organization=Graduate School of Mathematical Sciences, The University of Tokyo,addressline=3-8-1 Komaba, Meguro-ku, city=Tokyo, postcode=153-8914, country=Japan
@@ 1 Introduction
The Euler characteristic of crepant resolutions is one of the important geometric invariants in the study of the McKay correspondence. As a series of results connecting algebra with geometry, the McKay correspondence studies relations between algebraic properties of groups and geometric properties of the associated quotient singularities. Over the complex numbers, it was conjectured by Reid and proved by Batyrev via motivic integration ([1]), that for a finite group $G\subseteq\mathrm{SL}(n,\mathbb{C})$ and the associated quotient singularity $X:=\mathbb{C}^{n}/G$, assuming the existence of a crepant resolution $Y\to X$, the Euler characteristic $e(Y)$ equals the number of conjugacy classes $\mathrm{Conj}(G)$.
In this paper, we focus on the case where the base field has positive characteristic. Let $k$ be an algebraically closed field of prime characteristic $p>0$, and $G\subseteq\mathrm{SL}(n,k)$ be a finite group giving the quotient singularity $X:=\mathbb{A}^{n}_{k}/G$. For a smooth $k$-variety, its (topological) Euler characteristic is defined as the alternating sum of the Betti numbers given by the $l$-adic cohomology with compact support, where $l$ is a prime different from $p$.
It is natural to ask whether the equality between the Euler characteristic of crepant resolutions and the number of conjugacy classes still holds. If the order of $G$ is not divisible by $p$, then there is an affirmative answer known as the tame McKay correspondence (the reader can refer to [13], Section 5.2). Unfortunately, in the modular case, where #G\#G is divisible by $p$, the analog of Batyrev’s theorem has been disproved by constructing counterexamples, such as Chen, Du and Gao’s counterexample in the non-small case ([3]), Yamamoto’s construction for the quotient variety given by the symmetric group $S_{3}\subseteq\mathrm{SL}(3)$ in characteristic $3$ ([15]) and the author’s construction for the quotient variety given by the permutation action of the alternating group $A_{4}$ in characteristic $2$ ([5]).
On the other hand, there are also several known results implying that the analog of Batyrev’s theorem in positive characteristic holds for some specific modular quotient singularities. We list them below.
@@
    ([17], Corollary 6.21) Let $k$ be an algebraically closed field of characteristic $p>0$. Suppose that $G\subseteq\mathrm{SL}(n,k)$ is a finite group with no pseudo-reflections, such that $G\cong C_{p}$, where $C_{p}$ is the $p$-cyclic group. If the associated quotient singularity has a crepant resolution $Y\to X:=\mathbb{A}^{n}_{k}/G$, then $e(Y)=\#\mathrm{Conj}(C_{p})=p$.
@@
    ([15], Theorem 1.2) Let $k$ be an algebraically closed field of characteristic $3$. Suppose that $G\subseteq\mathrm{SL}(3,k)$ is a finite group with no pseudo-reflections, such that $G\cong H\rtimes C_{3}$, where $H$ is a non-modular abelian group. Then the associated quotient singularity has a crepant resolution $Y\to X:=\mathbb{A}^{3}_{k}/G$, and $e(Y)=\#\mathrm{Conj}(G)$.
@@
    ([4], Corollary 3.1) Let $k$ be an algebraically closed field of characteristic $2$. For any positive odd number $l$, let $\zeta_{l}$ be a primitive $l$-th root of unity in $k$. For integers $a_{1},a_{2},a_{3},a_{4}$, denote the diagonal matrix $\mathrm{diag}(\zeta_{l}^{a_{1}},\zeta_{l}^{a_{2}},\zeta_{l}^{a_{3}},\zeta_{l}^{a_{4}})$ by $\frac{1}{l}(a_{1},a_{2},a_{3},a_{4})$. Suppose that $G\subseteq\mathrm{SL}(4,k)$ has a semidirect product structure $G\cong H\rtimes C_{2}$, where $C_{2}$ is generated by the permutation $(12)(34)$ as an element of $\mathrm{SL}(4,k)$, and $H$ is one of the following:
    • (a)
@@
    Then $G$ has no pseudo-reflections, and the associated quotient singularity has a crepant resolution $Y\to X:=\mathbb{A}^{4}_{k}/G$, satisfying $e(Y)=\#\mathrm{Conj}(G)$.
In the results above, $G$ is always small (that is, $G$ has no pseudo-reflections) and has a semidirect product structure given by a non-modular abelian normal subgroup and a $p$-cyclic group. If $G\cong H\rtimes K$, then it is a classical idea to consider the decomposition $\mathbb{A}^{n}\xrightarrow{/H}\mathbb{A}^{n}/H\xrightarrow{/K}\mathbb{A}^{n}/G$ and the singularities given by $H$ and $K$ respectively. Under the assumption, $H$ is non-modular, and the singularity given by $K\cong C_{p}$ is studied by Theorem 1.1.1, with the same statement as the Batyrev’s theorem. In fact, Theorem 1.1.2 and Theorem 1.1.3 are parallel to Ito’s construction of crepant resolutions of trihedral singularities ([8]) and specific $4$-dimensional quotient singularities ([6]) over $\mathbb{C}$, with the same approach. In this paper, we show that for such small groups with semidirect product structures, the analogous statement of the Batyrev’s theorem always holds.
@@
Let $k=\overline{\mathbb{F}_{p}}$ be the algebraic closure of the finite fields of characteristic $p>0$, and $G\subseteq\mathrm{SL}(n,k)$ be a finite small group, such that $G$ has a non-modular abelian normal subgroup $H$ of index $p$. If $X:=\mathbb{A}^{n}_{k}/G$ has a crepant resolution $f:Y\to X$, then
$e(Y)=\#\mathrm{Conj}(G)=\#\mathrm{Ind}_{k}(G),$
where $\mathrm{Ind}_{k}(G)$ is the set of indecomposable $kG$-modules up to isomorphism.
Thus, Theorem 1.1.2 and Theorem 1.1.3 become examples of the main theorem. By this result, we also point out that the number of isomorphism classes of indecomposable $kG$-modules, which is not necessarily equal to $\mathrm{Conj}(G)$ in positive characteristic, is a candidate for the algebraic invariant of $G$ corresponding to the Euler characteristic of the crepant resolution. Indeed, we have the following conjecture.
@@
Let $k$ be an algebraically closed field of characteristic $p>0$, and $G\subseteq\mathrm{SL}(n,k)$ be a finite small group of finite representation type (that is, there are finitely many isomorphism classes of indecomposable $kG$-modules). Let $P$ be a $p$-Sylow subgroup of $G$. If both $X:=\mathbb{A}^{n}_{k}/G$ and $X^{\prime}:=\mathbb{A}^{n}_{k}/P$ have crepant resolutions $f:Y\to X,f^{\prime}:Y^{\prime}\to X^{\prime}$, then
$e(Y)=\#\mathrm{Ind}_{k}(G).$
Remark 1.4.
The assumption of the existence of a crepant resolution of $\mathbb{A}^{n}_{k}/P$ is necessary here. Otherwise, $\mathbb{A}_{k}^{6}/S_{3}$ in characteristic 33 gives a counterexample, where the symmetric group $S_{3}$ acts on $\mathbb{A}_{k}^{6}$ by the direct sum of two copies of the permutation representation. See Chapter 7.4 of [2] and Section 5.4 of [13] for more details.
This conjecture is true for all known cases, including the non-modular case, the main theorem and Yamamoto’s counterexamples in [15]. On the other hand, there are also examples where the group is of infinite representation type, such as the author’s construction in [5]. In the last section, two quotient singularities given by certain groups of infinite representation type are considered.
@@
Let $k$ be an algebraically closed field of characteristic $2$, $A_{4}$ and $C_{2}^{2}$ be the subgroups of $\mathrm{SL}(4,k)$ given by the permutation actions of the alternating group and its $2$-Sylow subgroup respectively. Denote the associated quotient singularities by $X_{1}:=\mathbb{A}^{4}/{C_{2}^{2}}$ and $X_{2}:=\mathbb{A}^{4}/A_{4}$. If there exists a crepant resolution $Y_{1}\to X_{1}$ ($resp.$ $Y_{2}\to X_{2}$), then $e(Y_{1})=6$ ($resp.$ $e(Y_{2})=10$).
This paper is organized as follows. In Section 2, we introduce the results and notations that are used for the proof of the main theorem, including the main tool - the mass formula version of the wild McKay correspondence. In Section 3, we prove the main theorem when $G$ is abelian. In Section 4, we prove the main theorem when $G$ is non-abelian. In Section 5, we provide two examples of computing the Euler characteristic of crepant resolutions for groups of infinite representation type, using the key idea of the proof of our main theorem.
@@ 2 Preliminaries
For the proof of our main theorem, the wild McKay correspondence as mass formulas is an important tool. In this version of the motivic wild McKay correspondence over finite fields, stringy motives are studied by their realization as the stringy point-counts $_{\mathrm{st}}$. Under some specific conditions, if there exists a crepant resolution, then some of their geometric properties can be studied via the stringy point-counts of the quotient varieties. For more details, the reader can refer to [19]. Here below we only list the results that are necessary for our proof.
@@
Let $K$ be a field and $G$ be a finite group. $M$ is called a GG-étale $K$-algebra, if $M$ is a finite étale $K$-algebra of degree #G\#G equipped with a $G$-action, such that $M^{G}=K$. Isomorphisms of GG-étale $K$-algebras are the isomorphisms of $K$-algebras that are $G$-equivariant. $G\text{-}\mathrm{\acute{E}t}(K)$ is the set consisting of the isomorphism classes of GG-étale $K$-algebras.
@@
Different actions of $G$ on the same $K$-algebra may result in different GG-étale $K$-algebras. However, conjugate endomorphisms of $G$ do not change the class in $G\text{-}\mathrm{\acute{E}t}(K)$. For each $M\in G\text{-}\mathrm{\acute{E}t}(K)$, there exists a subgroup $G_{M}\subseteq G$ (that is unique up to conjugation) such that $M\cong L_{M}^{\oplus[G:G_{M}]}$, where $L_{M}$ is a Galois extension of $K$ with its Galois group isomorphic to $G_{M}$.
Theorem 2.8 ([19], Proposition 8.5).
Let $k$ be a finite field, $\mathbb{F}_{q}/k$ be a field extension with order $q=p^{e}$, and $K=\mathbb{F}_{q}((t))$ be the local field of formal Laurent series. Let $G$ be a finite small subgroup of $\mathrm{SL}(n,k)$. Let $X:=\mathbb{A}^{n}/G$ be the associated quotient variety. Then
$_{\mathrm{st},\mathbb{F}_{q}}X=\sum_{M\in G\text{-}\mathrm{\acute{E}t}(K)}\frac{q^{n-\mathbf{v}_{V}(M)}}{\#C_{G}(G_{M})};$
@@
$Y(\mathbb{F}_{q})=\#_{\mathrm{st},\mathbb{F}_{q}}X=\sum_{M\in G\text{-}\mathrm{\acute{E}t}(K)}\frac{q^{n-\mathbf{v}_{V}(M)}}{\#C_{G}(G_{M})}.$
Here $Y(\mathbb{F}_{q})$ counts the number of $\mathbb{F}_{q}$-rational points of $Y$, $C_{G}(G_{M})$ is the centralizer of $G_{M}$ in $G$, and $\mathbf{v}_{V}(M)$ is the $\mathbf{v}$ function defined as follows.
Definition 2.9 ([19]; $\mathbf{v}$ function).
Let $k$, $\mathbb{F}_{q}$ and $K$ be the fields of characteristic $p$ as stated in Theorem 2.8. Let $\mathcal{O}_{K}:=\mathbb{F}_{q}[[t]]$ be the ring of integers. $\mathbf{v}$ function is defined for an arbitrary finite group $G$ with a representation $\rho:G\to\mathrm{GL}(n,k)\hookrightarrow\mathrm{GL}(n,\mathcal{O}_{K})$. Let $V=\mathbb{A}^{n}_{\mathcal{O}_{K}}$ be the representation space and $\mathcal{O}_{K}[x_{1},\dots,x_{n}]$ be its coordinate ring with the linear part $T=\sum\mathcal{O}_{K}x_{i}$. Then for any $M\in G\text{-}\mathrm{\acute{E}t}(K)$, $G$ has both an action on $\mathcal{O}_{M}$ via the definition of GG-étale $K$-algebras and an action on $T$ via the representation $\rho$. Denote by $\mathrm{Hom}_{\mathcal{O}_{K}}^{G}(T,\mathcal{O}_{M})$ the homomorphisms that commute with the actions of $G$. Then $\mathrm{Hom}_{\mathcal{O}_{K}}^{G}(T,\mathcal{O}_{M})$ is a free $\mathcal{O}_{K}$-module of rank $n$ ([18], Proposition 6.3), and
@@
$\mathbf{v}$ function is convertible and additive ([13], Lemma 3.4). In this paper, we will use the following facts: Even if $\rho$ is not faithful, it can be decomposed as $G\rightarrow G/H\rightarrow\mathrm{GL}(n,k)$ for some $H\subseteq G$, and $\mathbf{v}_{V}(M)=\mathbf{v}_{V_{G/H}}(M^{H})$; if $V$ has a direct decomposition as $V=\bigoplus_{i=1}^{s}V_{i}$, then $\mathbf{v}_{V}=\sum_{i=1}^{s}\mathbf{v}_{V_{i}}$.
@@
The value of the $\mathbf{v}$ function for $M\in G\text{-}\mathrm{\acute{E}t}(K)$ is determined by $L_{M}/K$ as a $G_{M}$-étale $K$-algebra. We simply write $\mathbf{v}(L)$ if $G_{M}$ is fixed.
Next, we give some notations and use them to rewrite the wild McKay correspondence such that the Euler characteristic of crepant resolutions can be studied.
@@
$F_{G}(q):=\sum_{[G^{\prime}]:G^{\prime}\subseteq G}\frac{1}{\#N_{G}(G^{\prime})}f_{G^{\prime}}.$
Here $N_{G}(G^{\prime})$ is the normalizer of $G^{\prime}$ in $G$, and $G^{\prime}]$ runs over the classes of subgroups of $G$ up to conjugation.
@@
Every element in $A(G)$ contains the information of a Galois extension $L/K$ with Galois group $\mathrm{Gal}(L/K)\cong G$ and the action of $G$ on it. For example, if $G\cong C_{l}$ for some prime $l$ ($l\neq p$) and the order of $k$ is large enough, then by Kummer theory, there are $l+1$ Galois extensions of $K=k((t))$ with $l$-cyclic Galois group up to isomorphism: one is unramified and the other $l$ extensions are totally ramified. On the other hand, for each extension, there are l1l-1 ways to map the given generator of $G$ to a nontrivial element in the Galois group, giving different actions up to isomorphism of GG-étale algebras. Hence $A(C_{l})$ has $l^{2}-1$ elements.
@@
Let $f$ be a Laurent polynomial with coefficients in $\mathbb{Q}$. We define $S(f):=f(1)$, which is equal to the sum of the coefficients of $f$. Therefore, $S$ is a homomorphism of $\mathbb{Q}$-algebras.
@@
Under the assumption of Theorem 2.8, if there exists a crepant resolution $Y$, then $F_{G}(T^{2})$ is a $\mathbb{Z}$-coefficient polynomial in $T$, and $e(Y)=S(F_{G})$.
@@
The first assertion is from the motivic wild McKay correspondence (construction of stringy motives and their realization as Poincaré polynomials of crepant resolutions via $\mathbb{L}\mapsto T^{2}$), and the second one follows; one can also check it by the functional equation in the Weil conjectures. The reader can refer to [17] for details. ∎
@@
Assume that the order of $k$ is large enough. If $G$ is non-modular, then $F_{G}$ is a polynomial in $q$ and $S(F_{G})=\#\mathrm{Conj}(G)$. If $G$ is also abelian, then $\sum_{G^{\prime}\subseteq G}f_{G^{\prime}}=(\#G)^{2}$.
@@
This is a corollary of the tame McKay correspondence ([13], Section 5.2). We also point out that $S(f_{G})=\#A(G)$ here. ∎
@@
If $G$ is isomorphic to a cyclic group of order $p$ and there exists a crepant resolution $Y\to X$ , then $F_{G}$ is a polynomial in $q$, such that $S(F_{G})=p$ and $S(f_{G})=p^{2}-1$.
@@
This is a corollary of the $p$-cyclic McKay correspondence ([17], Corollary 6.21). ∎
Then, we give the proof of the equality between the number of conjugacy classes and the number of isomorphism classes of indecomposable representations in the statement of our main theorem.
@@
Let $G$ be a finite group equipped with a semidirect product structure $H\rtimes C_{p}$, where $H$ is abelian and $C_{p}$ is a cyclic group of prime order $p$, such that $p\nmid\#H$. Let $C(G)$ be the center of $G$ and $k$ be a splitting field of $G$ of characteristic $p$ (equivalently, for any element $h\in H$, $k$ contains all of the $\mathrm{ord}(h)$-th primitive roots of unity in $\overline{k}$). Then
$\mathrm{Conj}(G)=\#\mathrm{Ind}_{k}(G)=\left\{\begin{array}[]{ll}\displaystyle p\#C(G)+\frac{\#H-\#C(G)}{p}&G\text{ is not abelian,}\\ \#G&G\text{ is abelian.}\end{array}\right$.